OpenStudy (elisaneedshelp):

Can a warm soul please help me with finding the value of 8^4/3 ? Pleassee

Mehek (mehek14):

what if I have a cold soul?

OpenStudy (elisaneedshelp):

\[8^{\frac{ 4 }{ 3 }}\] it be nice if you still help me

OpenStudy (elisaneedshelp):

the world is full of different races ethnicities and temperatures....

Mehek (mehek14):

idk how to do it :P

OpenStudy (owlcoffee):

There is a nice property of exponents that can help you here, it has the follwing form: \[A ^{\frac{ m }{ n }}=\sqrt[n]{A^m}\] This only means, that any exponent that is composed of a fraction, can be respresented as a radical, straight forward. So, in your excercise: \[8^{\frac{ 4 }{ 3 }}\] has the same form as the property, being m=4 and n=3 so, let´s transform it: \[8^{\frac{ 4 }{ 3 }}=\sqrt[3]{8^4}\] but 8^4 = 4096 so: \[\sqrt[3]{8^4}= \sqrt[3]{4096}\] and you can calculate that with your calculator.

OpenStudy (elisaneedshelp):

Thank you :)